OpenStudy (blondisaurus):

Solve by factoring. 4x^2+28x-32=0

OpenStudy (anonymous):

do it by middle term spilt

OpenStudy (anonymous):

4x^2 + 32x - 4x -32 = 0

OpenStudy (jiteshmeghwal9):

use the quadratic formula\[\LARGE{\color{red}{-b \pm \sqrt{b^2-4ac}\over2a}}\]here, a=4 b=28 c=-32\[\LARGE{-(28) \pm \sqrt{(28)^2-4(4)(-32)}\over2(4)}\]

OpenStudy (jiteshmeghwal9):

now solve this equation to gt ur answer @Blondisaurus :)

OpenStudy (jiteshmeghwal9):

u can alsp choose the method of @uzumakhi he is also right and his method is also easier :)

hartnn (hartnn):

@Blondisaurus did u get why 28x was split into 32x-4x ?

OpenStudy (anonymous):

Take out common factor of four

OpenStudy (anonymous):

Thanx @hartnn this is my next question to @Blondisaurus

OpenStudy (jiteshmeghwal9):

but for using the method of @uzumakhi u must learn about factorization :)

OpenStudy (anonymous):

i think @Blondisaurus is gone

OpenStudy (blondisaurus):

No im here. Just trying to make sense of it

OpenStudy (anonymous):

so it's clear or you are having any problem

OpenStudy (blondisaurus):

Still kinda unclear on how the answer ended up being 1,-8

OpenStudy (anonymous):

so do you understand middle term spilt

OpenStudy (blondisaurus):

No...UGH i feel retarded. >.<

OpenStudy (anonymous):

to know about middle term split you have to understand factorization

OpenStudy (anonymous):

to understand it go to following address http://www.teacherschoice.com.au/Maths_Library/Algebra/Alg_18.htm

OpenStudy (blondisaurus):

Okay I'll look over it later, right now i gotta get ready for class but thanks for going out of your way to help. Can see myself using this website alot this semester haha. Thanks again to everyone! :D

OpenStudy (anonymous):

ok good-luck